### v1

# Thermostats without conjugate points

Abstract.
We generalize Hopf’s theorem to thermostats: the total thermostat curvature of a thermostat without conjugate points is non-positive, and vanishes only if the thermostat curvature is identically zero. We further show that, if the thermostat curvature is zero, then the flow has no conjugate points, and the Green bundles collapse almost everywhere. Given a thermostat without conjugate points, we prove that the Green bundles are transversal everywhere if and only if it admits a dominated splitting. Finally, we provide an example showing that Hopf’s rigidity theorem on the $2$-torus cannot be extended to thermostats. It is also the first example of a thermostat with a dominated splitting which is not Anosov.
1. Introduction
Thermostats model the motion of a particle moving on a surface under the influence of a force acting orthogonally to the velocity. Unlike the special case of magnetic flows, these systems allow the force to depend on the particle’s velocity, yielding examples of dissipative flows that still preserve the initial kinetic energy. As such, they provide interesting models in non-equilibrium statistical mechanics, as studied by Gallavotti and Ruelle in [GR97, Gal99, Rue99].
Concretely, let $(M,g)$ be a closed oriented Riemannian surface, and let $\lambda\in\mathcal{C}^{\infty}(SM,\mathbb{R})$ be a smooth function on the unit tangent bundle $\pi:SM\to M$. A curve $\gamma:\mathbb{R}\rightarrow M$ is a thermostat geodesic if it satisfies the equation
$\nabla_{\dot{\gamma}}\dot{\gamma}=\lambda(\gamma,\dot{\gamma})J\dot{\gamma},$
where $\nabla$ is the Levi-Civita connection associated to $g$ and $J:TM\rightarrow TM$ is the complex structure on $M$ induced by the orientation. Since the speed of $\gamma$ remains constant, this equation determines a flow on $SM$ given by $\varphi_{t}(\gamma(0),\dot{\gamma}(0)):=(\gamma(t),\dot{\gamma}(t))$. Its infinitesimal generator is $F:=X+\lambda V$, where $X$ is the geodesic vector field and $V$ is the vertical vector field. The system $(M,g,\lambda)$ is called a thermostat.
The way in which $\lambda$ depends on the velocity can have drastic effects on key dynamical properties of the flow. This richness also means that thermostats can encode a wide range of dynamical systems. For instance:
  1. (a)
    Geodesic flows ($\lambda=0$). These are contact and reversible in the sense that the flip $(x,v)\mapsto(x,-v)$ on $SM$ conjugates $\varphi_{t}$ with $\varphi_{-t}$.
  2. (b)
    Magnetic flows ($\lambda$ does not depend on the velocity). These are volume-preserving, irreversible, and can behave very differently based on the initial kinetic energy.
  3. (c)
    Gaussian thermostats ($\lambda$ depends linearly on the velocity). These are reversible, not volume-preserving, and may not preserve an absolutely continuous measure. They correspond to geodesic flows of Weyl connections, which can have torsion.
  4. (d)
    Quasi-Fuchsian flows ($\lambda$ is the real part of a holomorphic quadratic differential). When these are Anosov, the weak stable and unstable bundles are smooth, yet the SRB measure is singular when $\lambda\neq 0$ and hence not algebraic.
  5. (e)
    Projective flows ($\lambda$ has linear and cubic velocity terms). These include the geodesic flows of torsion-free affine connections, up to parametrization.
  6. (f)
    Coupled vortex equations ($\lambda$ is the real part of a holomorphic differential of degree $m\geq 2$). These are related to the Ginzburg–Landau model for superconductors.
The goal of this paper is to explain how the no-conjugate-points condition relates to different notions of curvature, as well as characterize the dynamics of such thermostats. In doing so, we highlight both the concepts that generalize perfectly from the geodesic case, and the nuances that appear with greater dynamical complexity. This exercise not only sheds new light for thermostats, but also gives new results for geodesic flows.
As in the Riemannian setting, we define the exponential map $\exp^{\lambda}:TM\to M$ by
$\exp_{x}^{\lambda}(tv):=\pi(\varphi_{t}(x,v)),\qquad x\in M,\quad t\geq 0,\quad v\in S_{x}M.$ (1.1)
For every $x\in M$, $\exp_{x}^{\lambda}$ is a $\mathcal{C}^{1}$ map on $T_{x}M$ which is $\mathcal{C}^{\infty}$ on $T_{x}M\setminus\{0\}$ (see, for instance, the proof of [DPSU07, Lemma A.7]). We say that the thermostat in question has no conjugate points if $\exp_{x}^{\lambda}$ is a local diffeomorphism for all $x\in M$.
1.1. Conjugate points and thermostat curvature
Let $K_{g}$ denote the Gaussian curvature of $(M,g)$. In the geodesic case, it is straightforward to check that $K_{g}\leq 0$ implies $g$ has no conjugate points. The quantity that usually plays the role of Gaussian curvature for thermostats is the thermostat curvature of $(M,g,\lambda)$, given by
$\mathbb{K}:=\pi^{*}K_{g}-H(\lambda)+\lambda^{2}+F(V(\lambda)),$
where $H:=[V,X]$. Note this is a function on $SM$ instead of $M$. It turns out that this notion of thermostat curvature is implicitly making a choice of a gauge (see §2.1). Given a function $p:SM\to\mathbb{R}$ differentiable along the flow, define
$\kappa_{p}:=\pi^{*}K_{g}-H(\lambda)+\lambda^{2}+F(p)+p(p-V(\lambda)).$ (1.2)
Observe that $\mathbb{K}$ is nothing but $\kappa_{p}$ with $p=V(\lambda)$. The quantity $\kappa_{p}$ was first defined in [MP19] as a way of understanding the dynamics of the thermostat flow.
Our first result shows that $\kappa_{p}$ offers a useful criteria to check whether a thermostat has no conjugate points.
Theorem 1.1.
Let $(M,g,\lambda)$ be a thermostat. If $\kappa_{p}\leq 0$ for some function $p:SM\to\mathbb{R}$ which is differentiable along the flow, then there are no conjugate points.
In fact, the additional degree of freedom coming from the gauge $p$ allows us to completely characterize thermostats without conjugate points. This characterization is also new for geodesic flows.
Theorem 1.2.
A thermostat $(M,g,\lambda)$ has no conjugate points if and only if there exists a Borel measurable function $p:SM\to\mathbb{R}$, smooth along the flow, with $\kappa_{p}=0$.
Next, we give a generalization of Hopf’s rigidity result in [Hop48] to thermostats. Note that $\mu$ denotes the Liouville form on $SM$.
Theorem 1.3.
Let $(M,g,\lambda)$ be a thermostat without conjugate points. For any Borel measurable function $p:SM\to\mathbb{R}$ differentiable along the flow, we have
$\int_{SM}\kappa_{p}\mu\leq\int_{SM}(p-V(\lambda))^{2}\mu.$ (1.3)
Moreover, if equality holds, then $\mathbb{K}=0$.
As a consequence of (1.3) with $p=V(\lambda)$, Stokes’ theorem, and the Gauss-Bonnet theorem, we get
$2\pi\chi(M)+\int_{SM}(\lambda^{2}-(V(\lambda))^{2})\mu\leq 0,$ (1.4)
where $\chi(M)$ is the Euler characteristic of $M$. Furthermore, observe that the exponential map in (1.1) cannot yield a covering of the $2$-sphere for topological reasons, so any surface admitting a thermostat without conjugate points must have genus at least one.
Let us briefly focus on the case where $M$ is homeomorphic to the $2$-torus $\mathbb{T}^{2}$. In the geodesic case (i.e., $\lambda=0$), we recover the classical fact by Hopf that $g$ must be flat. In the magnetic case, i.e., $V(\lambda)=0$, the inequality (1.4) implies that $\lambda=0$, and then $g$ must also be flat (see also [AMRT24, Corollary C]). In some sense, these observations are telling us that the situation on $\mathbb{T}^{2}$ is very rigid when the flow is volume-preserving. If we also allow $\lambda$ to have a linear term with respect to the velocity, [AD14, Theorem 1.1] tells us that the magnetic component of $\lambda$ (i.e., the zeroth Fourier mode $\lambda_{0}$) must still identically vanish, and the metric $g$ must be conformally flat. The following no-go result shows that this rigidity does not apply to more general thermostats on $\mathbb{T}^{2}$.
Theorem 1.4.
For any Riemannian metric $g$ on $\mathbb{T}^{2}$, there exists $\lambda\in\mathcal{C}^{\infty}(S\mathbb{T}^{2},\mathbb{R})$ such that the thermostat $(\mathbb{T}^{2},g,\lambda)$ has no conjugate points. Moreover, the function $\lambda$ can always be chosen such that $\lambda_{0}\neq 0$.
1.2. Green bundles in the cotangent bundle
A key observation in the study of metrics without conjugate points is the existence of two flow-invariant bundles over the unit tangent bundle, known as Green bundles. The construction of these bundles was extended to the setting of convex Hamiltonians in [CI99]. However, these arguments do not directly carry over to the thermostat setting, as thermostats may be dissipative.
We take a new approach to understanding the Green bundles by working on the cotangent bundle $T^{*}(SM)$ as opposed to the tangent bundle $T(SM)$. The primary motivation is that, if we consider the induced dynamics, there is a smooth invariant subbundle $\Sigma\subset T^{*}(SM)$ which, in spirit, can replace the notion of a contact distribution in the geodesic case. Indeed, the symplectic lift of $\varphi_{t}$ to $T^{*}(SM)$, given by
$\tilde{\varphi}_{t}(v,\xi):=\left(\varphi_{t}(v),d\varphi_{t}^{-\top}(v)\xi\right),\qquad(v,\xi)\in T^{*}(SM),$ (1.5)
is the Hamiltonian flow of $\xi(F(v))$, so it preserves the characteristic set $\Sigma$ with fibers
$\Sigma(v):=\{\xi\in T_{v}^{*}SM\ |\ \xi(F(v))=0\}.$ (1.6)
When working on $T^{*}(SM)$, it is natural to introduce a moving coframe. The vector fields $\{X,H,V\}$ form an orthonormal frame for $T(SM)$ with respect to the Sasaki metric (the natural lift of $g$ to $SM$). We can then consider the corresponding dual frame $\{\alpha,\beta,\psi\}$ for $T^{*}(SM)$. The cohorizontal distribution is defined as $\mathbb{H}^{*}:=\mathbb{R}\beta$, whereas the covertical is $\mathbb{V}^{*}:=\mathbb{R}\psi$. In the geodesic case, we have $\Sigma=\mathbb{H}^{*}\oplus\mathbb{V}^{*}$. For a thermostat, we introduce $\psi_{\lambda}:=\psi-\lambda\alpha$ and $\mathbb{V}_{\lambda}^{*}:=\mathbb{R}\psi_{\lambda}$ so that $\Sigma=\mathbb{H}^{*}\oplus\mathbb{V}^{*}_{\lambda}$.
Figure 1. The relevant subbundles in cotangent space.
In this new setting, we construct the Green bundles on orbits without conjugate points. Note that these are usually defined as subbundles of $T(SM)$ instead of $T^{*}(SM)$.
Theorem 1.5.
Let $(M,g,\lambda)$ be a thermostat. For a given $v\in SM$, either:
  1. (a)
    There exist $t_{0}\in\mathbb{R}$ and $t\neq 0$ such that
    $d\varphi_{t}^{-\top}(\mathbb{H}^{*}(\varphi_{t_{0}}(v))=\mathbb{H}^{*}(\varphi_{t_{0}+t}(v)).$ (1.7)
    This happens if and only if $t\mapsto\pi(\varphi_{t}(v))$ contains conjugate points on $M$.
  2. (b)
    There exist two invariant subbundles $G_{s/u}^{*}\subset\Sigma$ along the orbit of $v$ given by
    $G^{*}_{s}(v):$ $=\lim_{t\to\infty}d\varphi_{-t}^{-\top}(\mathbb{H}^{*}(\varphi_{t}(v)),$ (1.8)
    $G^{*}_{u}(v):$ $=\lim_{t\to\infty}d\varphi_{t}^{-\top}(\mathbb{H}^{*}(\varphi_{-t}(v)).$
    They satisfy the transversality condition
    $G^{*}_{s}(v)\cap\mathbb{H}^{*}(v)=\{0\}=G^{*}_{u}(v)\cap\mathbb{H}^{*}(v).$ (1.9)
In [Ebe73, Theorem 3.2], Eberlein characterized Anosov geodesic flows as geodesic flows without conjugate points which have transversal Green bundles. In [CI99, Theorem C], a similar statement was given for convex Hamiltonians (which might be non-contact, but are still volume-preserving). Recall that a flow $(\varphi_{t})_{t\in\mathbb{R}}$ is Anosov if there is a flow-invariant splitting $T(SM)=\mathbb{R}F\oplus E^{s}\oplus E^{u}$ as well as uniform constants $C\geq 1$ and $10<\rho<$ 1 such that for all $t\geq 0$ we have
$\|d\varphi_{t}|_{E^{s}}\|\leq C\rho^{t},\qquad\|d\varphi_{-t}|_{E^{u}}\|\leq C\rho^{t}.$ (1.10)
As we will see, due to the lack of volume preservation, the natural extension of these results to thermostats applies to dominated splittings instead. The flow $(\varphi_{t})_{t\in\mathbb{R}}$ is said to admit a dominated splitting if the estimates (1.10) are relaxed to
$\|\left.d\varphi_{t}\right|_{E^{s}(v)}\|\|\left.d\varphi_{-t}\right|_{E^{u}(\varphi_{t}(v))}\|\leq C\rho^{t}$ (1.11)
for all $t\geq 0$. This can be regarded as a projective form of the Anosov property. For geodesic and magnetic flows, the Anosov property is equivalent to having a dominated splitting due to volume preservation.
The thermostat version of Eberlein’s result is hence the following.
Theorem 1.6.
Let $(M,g,\lambda)$ be a thermostat without conjugate points. It admits a dominated splitting if and only if $G^{*}_{s}(v)\cap G^{*}_{u}(v)=\{0\}$ for all $v\in SM$.
Remark 1.7.
In establishing the result for geodesic flows, Eberlein uses Klingenberg [Kli74] to argue that an Anosov geodesic flow must be without conjugate points. In absence of such a result for thermostats admitting dominated splittings, we assume that the thermostat is without conjugate points.
If the thermostat has no conjugate points and admits a dominated splitting, then $G^{*}_{s}=(E^{s})^{*}$ and $G^{*}_{u}=(E^{u})^{*}$, where the dual bundles are defined by
$(E^{s})^{\ast}(\mathbb{R}F\oplus E^{s})=0=(E^{u})^{\ast}(\mathbb{R}F\oplus E^{u}).$ (1.12)
One can check that we have similar estimates to (1.11) for $(E^{s})^{*}$ and $(E^{u})^{*}$, with $d\varphi_{t}$ replaced by $d\varphi_{t}^{-\top}$.
In order to prove Theorem 1.6, we use the following characterization of thermostats with transversal Green bundles. Interestingly, even in the well-studied geodesic case, it provides a new partial converse to [Ano67]: even though the Anosov property does not imply negative Gaussian curvature, it does tell us that the thermostat curvature in an appropriate gauge is negative everywhere.
Theorem 1.8.
If a thermostat $(M,g,\lambda)$ without conjugate points has transversal Green bundles, there is a continuous $p:SM\to\mathbb{R}$, differentiable along the flow, with $\kappa_{p}<0$.
Finally, we explore the other extreme; namely, when the Green bundles collapse to a line everywhere instead of being transversal. For geodesic flows without conjugate points, a conjecture of Freire and Mañé can be rephrased as stating that the Green bundles collapse if and only if the metric is flat [FMn82]. Note that the Green bundles collapsing implies that the fundamental group grows subexponentially, and thus, if $M$ is a surface, we have that it must be a sphere or a torus. The conjecture for surfaces then follows by [Hop48].
A natural question is whether this extends to the thermostat setting with the thermostat curvature in place of the Gaussian curvature. As we will see in Proposition 4.1, this is not the case. In spite of this, there is still a connection between the Green bundles collapsing to a line, the thermostat curvature $\mathbb{K}$, and a quantity which we refer to as the damped thermostat curvature
$\tilde{\kappa}:=\pi^{*}K_{g}-H(\lambda)+\lambda^{2}+\frac{F(V(\lambda))}{2}-\frac{(V(\lambda))^{2}}{4}.$ (1.13)
Note that $\tilde{\kappa}$ is simply $\kappa_{p}$ with the particular choice of $p=V(\lambda)/2$.
Theorem 1.9.
Let $(M,g,\lambda)$ be a thermostat.
  1. (a)
    If $\mathbb{K}=0$, then the flow has no conjugate points and $G^{*}_{s}=G^{*}_{u}=\mathbb{R}(\psi_{\lambda}-V(\lambda)\beta)$ $\mu$-almost everywhere. Moreover, if $V(\lambda)=0$, then $G^{*}_{s}=G^{*}_{u}=\mathbb{R}\psi_{\lambda}$ everywhere.
  2. (b)
    If $\tilde{\kappa}\leq 0$, then for any invariant Borel measure $\nu$ on $SM$ we have $\tilde{\kappa}=0$ $\nu$-almost everywhere if and only if $G^{*}_{s}=G^{*}_{u}$ $\nu$-almost everywhere.
(a) Conjugate points
(b) Dominated splitting
(c) $\mathbb{K}=0$
Figure 2. The lifted dynamics on $\Sigma$.
In Section 4, we will show that Theorem 1.9 (a) is optimal: when $V(\lambda)\neq 0$, it is possible to have $\mathbb{K}=0$ yet $G^{*}_{s}(v)\cap G^{*}_{u}(v)=\{0\}$ for some $v\in SM$. In particular, this implies that the conjecture of Freire and Mañé does not extend to the setting of thermostats with the thermostat curvature in place of the Gaussian curvature. On the other hand, observe that if $V(\lambda)=0$ (i.e., the system is magnetic), then $\mathbb{K}=\tilde{\kappa}=\kappa_{0}$, and $\mu$ becomes an invariant measure for the thermostat flow. Thus, Theorem 1.9 implies the following.
Corollary 1.10.
Let $(M,g,\lambda)$ be a magnetic system with $\kappa_{0}\leq 0$. We have $\kappa_{0}=0$ if and only if $G_{s}^{*}=G_{u}^{*}$ everywhere.
Combining the above results together with [MP19, Proposition 3.5, Theorem 3.7], we have the following picture:
[…]
This should be contrasted with the following, which summarizes what was previously known in the geodesic setting:
${K_{g}<0}$${K_{g}\leq 0}$Anosov no conjugate points + transversal Green bundles no conjugate points.
1.3. Remaining questions
As noted in Remark 1.7, it is not clear whether one needs the assumption that the thermostat is without conjugate points in Theorem 1.6. Given an Anosov flow on $SM$, [Ghy84, Theorem A] tells us that it is topologically orbit equivalent to the geodesic flow of any metric of constant negative Gaussian curvature on $M$. Therefore these flows are transitive, and their non-wandering set is all of $SM$. This property ends up being critical in proving that there are no conjugate points. In contrast, as we will see with concrete examples in Section 4, thermostats with dominated splittings can have non-trivial wandering sets.
Question. Can a thermostat with a dominated splitting have conjugate points?
One possible approach to this problem is to try to understand thermostats on the $2$-sphere. As pointed out above, it is easy to see using the exponential map that every such thermostat must have conjugate points. If one can construct an example which admits a dominated splitting, then this would show that the dominated splitting assumption is not enough. However, it is not even clear whether there can be an arbitrary flow with a dominated splitting on the unit tangent bundle of the $2$-sphere; the work of Arroyo and Rodriguez Hertz [ARH03] gives some insight into the problem, but it does not seem to be enough to rule out such examples.
Question. Is there a flow on the unit tangent bundle of the $2$-sphere which admits a dominated splitting?
Another possible approach to this problem is to try to understand whether all dominated splittings give rise to hyperbolic behavior. In Section 4, we give explicit examples of thermostats on the $2$-torus which admit dominated splittings and are not Anosov, showing that this is not the case, and answering a question of Mettler and Paternain [MP19] about the existence of such systems. These examples critically rely on the fact that the surface is a $2$-torus, and thus it may be possible that a thermostat with a dominated splitting will always be Anosov if the surface is not the $2$-torus. In light of the previous question, one can try exploring the following.
Question. Is there a thermostat on a surface of genus at least two which admits a dominated splitting and is not Anosov?
Next, we suspect that the assumption $\kappa_{0}\leq 0$ is not required in Corollary 1.10, and hence the weaker conjecture does extend to the setting of magnetic systems on surfaces. Provided the function $\lambda$ is sufficiently nice (i.e., if the Mañé critical value is less than $1/2$), one can use Theorem 1.3 and [BP02b, Theorem D and Proposition 5.4] to deduce the result. It is not immediately obvious how to show that $\kappa_{0}=0$ if one only assumes that the topological entropy is zero and the genus of $M$ is at least two.
Question. Let $(M,g,\lambda)$ be a magnetic system without conjugate points. Does $G_{s}^{*}=G_{u}^{*}$ everywhere imply $\kappa_{0}=0$?
Based on Theorem 1.9 (b), it is possible that the damped thermostat curvature is better equipped for detecting the topological entropy of the thermostat flow. Even in the setting where $\tilde{\kappa}\leq 0$, though, it is not clear whether $\tilde{\kappa}=0$ everywhere is equivalent to the Green bundles collapsing everywhere.
Question. Let $(M,g,\lambda)$ be a thermostat. Is $\tilde{\kappa}=0$ if and only if $G_{s}^{*}=G_{u}^{*}$ everywhere?
1.4. Organization of the paper
In Section 2, we study the relationship between conjugate points, Green bundles, and thermostat curvature. We describe the lifted dynamics of the thermostat on the characteristic set in §2.1. In §2.2, we explain how our point of view is equivalent to the one using cocycles. The lifted dynamics give us an interpretation of the no-conjugate-points condition in terms of the twist property of the cohorizontal bundle in §2.3, unlocking Theorem 1.1. Next, in §2.4, we construct the Green bundles, and prove Theorems 1.2, 1.3, and 1.5. We finish the section by studying the relationship between the Lyapunov exponents of the flow and the Green bundles in §2.5, giving us the tools to prove Theorem 1.9.
In Section 3, we explore the relationship between dominated splittings and transversal Green bundles. We first show that transversal Green bundles must be continuous and then use this property to prove Theorems 1.6 and 1.8.
Finally, in Section §4, we present a family of examples of thermostats on $\mathbb{T}^{2}$ without conjugate points which admit dominated splittings (yet are not Anosov), proving Theorem 1.4.
Acknowledgements
We are grateful to Gabriel Paternain for feedback on early drafts and helpful discussions.
2. Conjugate points and Green bundles
In what follows, $(M,g)$ is a closed oriented Riemannian surface, and we take an arbitrary $\lambda\in\mathcal{C}^{\infty}(SM,\mathbb{R})$.
2.1. Dynamics on the characteristic set
Recall that $\{\alpha,\beta,\psi\}$ is the moving coframe for $SM$ dual to the orthonormal frame $\{X,H,V\}$ under the Sasaki metric. By using the commutation formulas
$[V,X]=H,\qquad[V,H]=-X,\qquad[X,H]=\pi^{*}(K_{g})V,$
we can derive the structure equations
$d\alpha=\psi\wedge\beta,\qquad d\beta=-\psi\wedge\alpha,\qquad d\psi=-\pi^{*}(K_{g})\alpha\wedge\beta.$
We will use an adapted coframe $\{\alpha,\beta,\psi_{\lambda}\}$, where $\psi_{\lambda}:=\psi-\lambda\alpha$, as it gives us nice coordinates on $\Sigma$. By writing $\mathbb{H}^{*}:=\mathbb{R}\beta$ and $\mathbb{V}_{\lambda}^{*}:=\mathbb{R}\psi_{\lambda}$, we get $\Sigma=\mathbb{H}^{*}\oplus\mathbb{V}^{*}_{\lambda}$. Combining the previous structure equations with Cartan’s formula, we obtain
$\mathcal{L}_{F}\alpha=\lambda\beta,\qquad\mathcal{L}_{F}\beta=\psi_{\lambda},\qquad\mathcal{L}_{F}\psi_{\lambda}=-\kappa_{0}\beta+V(\lambda)\psi_{\lambda},$
where $\kappa_{0}$ is defined in (1.2) (with $p=0$).
For any $p:SM\to\mathbb{R}$ differentiable along the flow, it will also be useful to define $\phi_{p}:=\psi_{\lambda}-p\beta$ so that $\{\alpha,\beta,\phi_{p}\}$ is an alternative coframe satisfying $\Sigma=\mathbb{H}^{*}\oplus\mathbb{R}\phi_{p}$. This is nothing but a change of coordinates on $\Sigma$. We now have
$\mathcal{L}_{F}\beta=p\beta+\phi_{p},\qquad\mathcal{L}_{F}\phi_{p}=-\kappa_{p}\beta+(V(\lambda)-p)\phi_{p}.$ (2.1)
Figure 3. The bases $\{\beta,\psi_{\lambda}\}$ and $\{\beta,\phi_{p}\}$ for $\Sigma$.
To each $\xi\in\Sigma(v)$, we associate the functions $x,y\in\mathcal{C}^{\infty}(\mathbb{R})$ characterized by
$d\varphi_{t}^{-\top}(v)\xi=x(t)\beta+y(t)\phi_{p}.$ (2.2)
They capture all the information of the lifted dynamics in $\Sigma$.
Lemma 2.1.
Let $\xi\in\Sigma(v)$. Along the orbit of $v$, we have the pair of equations
$\begin{cases}\dot{x}+px-\kappa_{p}y=0,\\ \dot{y}+x+(V(\lambda)-p)y=0.\end{cases}$ (2.3)
In particular, the $y$ component always satisfies the Jacobi equation
$\ddot{y}+V(\lambda)\dot{y}+\mathbb{K}y=0.$ (2.4)
Proof.
By the definition of the inverse transpose $d\varphi_{t}^{-\top}(v)$, we have
$\eta(t):=d\varphi_{t}^{-\top}(v)\xi\in T^{*}_{\varphi_{t}(v)}(SM)$
if and only if
$\xi=\eta(t)\circ d\varphi_{t}(v).$
Therefore, we may write
$\xi=x(t)\beta\circ d\varphi_{t}(v)+y(t)\phi_{p}\circ d\varphi_{t}(v)=x(t)\varphi_{t}^{*}\beta+y(t)\varphi_{t}^{*}\phi_{p}.$
Differentiating this identity with respect to $t$ and using the definition of the Lie derivative $\mathcal{L}_{F}$, we obtain
$0=\dot{x}\varphi_{t}^{*}\beta+x\varphi_{t}^{*}(\mathcal{L}_{F}\beta)+\dot{y}\varphi_{t}^{*}\phi_{p}+y\varphi_{t}^{*}(\mathcal{L}_{F}\phi_{p}).$
Using the equations (2.1), we see that
$0=(\dot{x}+px-\kappa_{p}y)\varphi_{t}^{*}\beta+(\dot{y}+x+(V(\lambda)-p)y)\varphi_{t}^{*}\phi_{p}.$
Since $\{\beta,\phi_{p}\}$ are linearly independent, we get the pair of equations (2.3), as desired. ∎
Remark 2.2.
Observe that this implies that $x$ is completely determined by $y$.
For each $v\in SM$, it will also be useful to introduce the damping function
[…]
Indeed, it allows us to associate to each $\xi\in\Sigma(v)$ a new function $z\in\mathcal{C}^{\infty}(\mathbb{R})$ given by
$z(t):=\dfrac{y(t)}{m(t)},$
which we think of as a damped $y$ component.
Lemma 2.3.
For each $\xi\in\Sigma(v)$, the $z$ component is a solution of the Jacobi equation
$\ddot{z}+\tilde{\kappa}z=0$ (2.6)
along the orbit of $v$, where $\tilde{\kappa}$ is the damped thermostat curvature defined in (1.13).
Proof.
First, we use the fact that $y=mz$ to get
$\dot{y}=-\dfrac{V(\lambda)}{2}y+m\dot{z}.$
Taking a second derivative, we obtain
$\ddot{y}$ $=-\dfrac{F(V(\lambda))}{2}y-\dfrac{V(\lambda)}{2}\left(-\dfrac{V(\lambda)}{2}y+m\dot{z}\right)-\dfrac{V(\lambda)}{2}m\dot{z}+m\ddot{z}$ (2.7)
$=m\left(-\dfrac{F(V(\lambda))}{2}z+\dfrac{(V(\lambda))^{2}}{4}z-V(\lambda)\dot{z}+\ddot{z}\right).$
On the other hand, the Jacobi equation (2.4) yields
$\ddot{y}$ $=-V(\lambda)\dot{y}-(\kappa+F(V(\lambda))y$
$=m\left(\dfrac{(V(\lambda))^{2}}{2}z-V(\lambda)\dot{z}-(\kappa+F(V(\lambda))z\right).$
Setting them equal to each other then gives us the claim since $m$ is nowhere zero. ∎
One of the advantages of studying the damped component $z$ instead of $y$ is that, thanks to Lemma 2.3, we will be able to use the following general result.
Lemma 2.4.
Let $k\in\mathcal{C}^{\infty}(\mathbb{R})$ be such that any non-trivial solution $z$ to the equation
$\ddot{z}+kz=0$ (2.8)
vanishes at most once. If such $z$ vanishes once, then $|z(t)|$ is unbounded as $t\to\pm\infty$.
Proof.
By normalizing, it suffices to consider the case where $z(0)=0$ and $\dot{z}(0)=1$. For each $t_{0}\neq 0$, we know thanks to the one-time vanishing property and the homogeneity of equation (2.8) that there exists a unique solution $z_{t_{0}}$ with $z_{t_{0}}(0)=1$ and $z_{t_{0}}(t_{0})=0.$
We claim that
$z_{t_{0}}=z_{-1}+\dfrac{z_{t_{0}}(-1)}{z(-1)}z.$
Indeed, since both sides satisfy (2.8) and agree at $t=-1$ and $t=0$, the one-time vanishing property tells us that they must agree for all $t\in\mathbb{R}$.
Differentiating with respect to $t$ and setting $t=0$, we obtain
$\dot{z}_{t_{0}}(0)=\dot{z}_{-1}(0)+\dfrac{z_{t_{0}}(-1)}{z(-1)}.$
Now let $t_{0}>0$. Since $z_{t_{0}}(-1)>0$ and $z(-1)<0$, we notice that $t_{0}\mapsto\dot{z}_{t_{0}}(0)$ is bounded above as $t_{0}\to\infty$. For $t>0$, note that the function
$t\mapsto z(t)\int_{t}^{t_{0}}\dfrac{d\tau}{(z(\tau))^{2}}$
satisfies (2.8). Moreover, using a Taylor expansion, we can see that it tends to $1$ as $t\to 0$. Since it also agrees with $z_{t_{0}}$ at $t=t_{0}$, the one-time vanishing property gives us
$z_{t_{0}}(t)=z(t)\int_{t}^{t_{0}}\dfrac{d\tau}{(z(\tau))^{2}}$ (2.9)
for all $t\in\mathbb{R}$. It follows that, for $t_{0}>t_{0}^{\prime}>0$, we have
$\dot{z}_{t_{0}}(0)-\dot{z}_{t_{0}^{\prime}}(0)=\int_{t_{0}^{\prime}}^{t_{0}}\dfrac{d\tau}{(z(\tau))^{2}}>0,$ (2.10)
so $t_{0}\mapsto\dot{z}_{t_{0}}(0)$ is monotone increasing as $t_{0}\to\infty$. Combined with the previous upper bound, this implies that $t_{0}\mapsto\dot{z}_{t_{0}}(0)$ converges, so we may take the limit $t_{0}\to\infty$ in (2.10) to obtain a convergent integral on the right-hand side. We conclude that $|z(t)|$ is unbounded as $t\to\infty$. The same argument works for $t\to-\infty$ if we take $t_{0}<0$. ∎
2.2. Cocyles
Using the moving coframe $\{\beta,\psi_{\lambda}\}$, we get an identification $\Sigma\cong M\times\mathbb{R}^{2}$. Therefore, for each $t\in\mathbb{R}$, we obtain a unique map $\Psi_{t}:SM\to\text{GL}(n,\mathbb{R})$ defined by
$\tilde{\varphi}_{t}(v,\xi)=(\varphi_{t}(v),\Psi_{t}(v)\xi)$
for all $(v,\xi)\in\Sigma\cong M\times\mathbb{R}^{2}$. The map $\Psi:SM\times\mathbb{R}\to\text{GL}(2,\mathbb{R})$ satisfies the cocycle property over the flow $\varphi_{t}$:
$\Psi_{t+s}(v)=\Psi_{s}(\varphi_{t}(v))\Psi_{t}(v).$
This is simply a different point of view of §2.1, with the explicit relationship given by
$\Psi_{t}(v):\begin{pmatrix}x(0)\\ y(0)\end{pmatrix}\mapsto\begin{pmatrix}x(t)\\ y(t)\end{pmatrix},$
where the $x$ and $y$ components satisfy the differential equations (2.3). If we denote by $\Phi:SM\to\mathfrak{gl}(2,\mathbb{R})$ the infinitesimal generator of $\Psi_{t}$, namely,
$\Phi(v):=\frac{d}{dt}\Big{|}_{t=0}\Psi_{t}(v),$
then equations (2.3) (with $p=0$) allow us to write
$\Phi=\begin{pmatrix}0&\kappa\\ -1&-V(\lambda)\end{pmatrix}.$
Of course, we could have made a different choice of coordinates on $\Sigma$. This is represented by a gauge, i.e., a smooth map $P:SM\to\text{GL}(2,\mathbb{R})$, which gives rise to a new cocyle over the flow $\varphi_{t}$ by conjugation:
$\tilde{\Psi}_{t}(v)=P^{-1}(\varphi_{t}(v))\Psi_{t}(v)P(v).$
One can check that the new infinitesimal generator $\tilde{\Phi}$ is related to $\Phi$ by
$\tilde{\Phi}=P^{-1}\Phi P+P^{-1}FP.$
Under this lens, our previous choice of basis $\{\beta,\phi_{p}\}$ corresponds to the gauge
$P=\begin{pmatrix}1&0\\ -p&1\end{pmatrix},$
and the infinitesimal generator of the new cocyle $\tilde{\Psi}_{t}$ becomes
$\tilde{\Phi}=\begin{pmatrix}-p&\kappa_{p}\\ -1&p-V(\lambda)\end{pmatrix}.$
Note in particular that, for $p=V(\lambda)/2$, we may write
$\tilde{\Phi}=-\frac{V(\lambda)}{2}\text{Id}+\begin{pmatrix}0&\tilde{\kappa}\\ -1&0\end{pmatrix},$ (2.11)
so that
$\tilde{\Psi}_{t}(v)=m(t)\Gamma_{t}(v),$ (2.12)
where $m$ is defined in (2.5), and $\Gamma_{t}$ is the cocycle generated by the last matrix in (2.11). In light of Lemma 2.3, we have
$\Gamma_{t}(v):\begin{pmatrix}x(0)\\ z(0)\end{pmatrix}\mapsto\begin{pmatrix}x(t)\\ z(t)\end{pmatrix}.$
Note that the infinitesimal generator of $\Gamma_{t}$ has trace zero, so $\Gamma_{t}:SM\to\text{SL}(2,\mathbb{R})$. This is essentially the same cocyle as the one we would get from a geodesic flow with Gaussian curvature $\tilde{\kappa}$; however, note that $\tilde{\kappa}$ is a function on $SM$ instead of $M$. By picking the right gauge and damping the $y$ component in the cotangent bundle, we are hence reducing the problem to something that resembles the geodesic case.
2.3. Conjugate points
Let $x_{0},x_{1}\in M$ be a pair of points such that there exists a thermostat geodesic $\gamma:[0,T]\rightarrow M$ with $\gamma(0)=x_{0}$ and $\gamma(T)=x_{1}$. We say that $x_{0}$ and $x_{1}$ are conjugate along $\gamma$ if $\exp_{x_{0}}^{\lambda}$ is singular at $T\dot{\gamma}(0)$. With this in mind, we observe the following.
Lemma 2.5.
Let $\gamma:[0,T]\rightarrow M$ be a thermostat geodesic with endpoints $\gamma(0)=x_{0}$ and $\gamma(T)=x_{1}$. The points $x_{0}$ and $x_{1}$ are conjugate along $\gamma$ if and only if there is a non-trivial solution $y$ to (2.4) satisfying $y(0)=y(T)=0$.
Proof.
Since the function $m$ defined in (2.5) is nowhere zero, we have $y(t)=0$ if and only if $z(t)=0$. We can thus conclude by applying Lemma 2.3 and [AD14, Theorem 4.3]. Indeed, while their Jacobi equation for $y$ is different from (2.4) because they are working in the tangent bundle, the authors show in [AD14, Section 5] that a change of variables puts their Jacobi equation in the same normal form as (2.6). ∎
Corollary 2.6.
A thermostat has no conjugate points if and only if there are no non-trivial solutions to (2.4) (or (2.6)) which vanish at two points.
Remark 2.7.
There is a nice geometric interpretation of what is going on in the cotangent bundle. Note that having $y(t)=0$ is equivalent to $d\varphi_{t}^{-\top}(v)(\xi)\in\mathbb{H}^{*}(\varphi_{t}(v))$, so we get part (a) of Theorem 1.5. See Figure 2 (a).
Armed with this perspective on conjugate points, we prove Theorem 1.1.
Proof of Theorem 1.1.
Let $z$ be a non-trivial solution to (2.6) and define
$w:=\left(\dot{z}-\left(p-\dfrac{V(\lambda)}{2}\right)z\right)z.$
By Lemma 2.3, we have
$\dot{w}$ $=\dot{z}^{2}-2\left(p-\dfrac{V(\lambda)}{2}\right)\dot{z}z-\left(\tilde{\kappa}+F\left(p-\dfrac{V(\lambda)}{2}\right)\right)z^{2}$
$=\left(\dot{z}-\left(p-\dfrac{V(\lambda)}{2}\right)z\right)^{2}-\left(\tilde{\kappa}+F\left(p-\dfrac{V(\lambda)}{2}\right)+\left(p-\dfrac{V(\lambda)}{2}\right)^{2}\right)z^{2}$
$=\left(\dot{z}-\left(p-\dfrac{V(\lambda)}{2}\right)z\right)^{2}-\kappa_{p}z^{2}.$
Since $\kappa_{p}\leq 0$ by assumption, we get $\dot{w}\geq 0,$ so the function $w$ is non-decreasing. Suppose for contradiction, using Corollary 2.6, that $z$ vanishes multiple times. Note that, because of (2.6), if $z$ vanishes on an interval, then we must have $z=0$ everywhere; thus, we may assume that $z$ vanishes on a discrete set. Let $t_{0}\in\mathbb{R}$ be such that $z(t_{0})=0$, and let $t_{1}:=\inf\{t>t_{0}\ |\ z(t)=0\}$. If $t_{1}=t_{0}$, then we have an infinite sequence $t_{n}\to t_{0}^{+}$ such that $z(t_{n})=0$. By the mean value theorem, we get a sequence $t_{n}^{\prime}\to t_{0}^{+}$ with $\dot{z}(t_{n}^{\prime})=0$ and hence $\dot{z}(t_{0})=z(t_{0})=0$ by continuity, forcing $z$ to be zero everywhere. Thus, we must have $t_{1}>t_{0}$.
By construction, $z$ does not vanish on the interval $(t_{0},t_{1})$. Since $w$ is non-decreasing and $w(t_{0})=w(t_{1})=0$, $w$ must vanish on this interval. But then $z$ solves the first order differential equation $\dot{z}=(p-V(\lambda)/2)z$ on $(t_{0},t_{1})$ with $z(t_{0})=z(t_{1})=0$; it is easy to see that this implies that $z=0$ on the interval, which gives a contradiction. ∎
2.4. Green bundles
Next, we want to show that being without conjugate points implies that the bundles $d\varphi_{t}^{-\top}(\mathbb{H}^{*}(\varphi_{-t}(v)))$ converge as $t\rightarrow\pm\infty$. In this paper, when we talk about convergence of bundles, we mean it in the sense that the bundles converge in the topology of the $4$-dimensional manifold $\Lambda(SM)$ obtained by projectivizing $\Sigma(v)$ for all $v\in SM$. In other words, $\Lambda(SM)$ is the bundle over $SM$ whose fiber over $v$ consists of all $1$-dimensional subspaces of $\Sigma(v)$. Note that $\mathbb{H}^{*}$ is a section of this bundle, and that $\tilde{\varphi}_{t}$ on $T^{*}(SM)$ naturally induces a flow on $\Lambda(SM)$.
Proof of Theorem 1.5.
Fix $v\in SM$. We have already shown part (a), so in what follows we may assume that $t\mapsto\pi(\varphi_{t}(v))$ contains no conjugate points on $M$. Equivalently, any solution $z$ to (2.6) vanishes at most once.
Let $z_{t_{0}}$ be as in the proof of Lemma 2.4. Using Remark 2.2, we see that $z_{t_{0}}$ uniquely determines a point $\xi_{t_{0}}\in\Sigma(v)$ so that $\mathbb{R}\xi_{t_{0}}=d\varphi_{t_{0}}^{-\top}(\mathbb{H}^{*}(\varphi_{-t_{0}}(v)))$. Note that the proof of Lemma 2.4 shows that $\dot{z}_{t_{0}}(0)$ converges as $t_{0}\rightarrow\pm\infty$. Using the continuous dependence of solutions to (2.6) on initial conditions, we have that the functions $z_{t_{0}}$ converges as $t_{0}\rightarrow\pm\infty$ to solutions $z_{\pm\infty}$ of (2.6) whose corresponding points $\xi_{\pm\infty}\in\Sigma(v)$ must span $\lim_{t_{0}\rightarrow\pm\infty}d\varphi_{t_{0}}^{-\top}(\mathbb{H}^{*}(\varphi_{-t_{0}}(v)))$, and the result follows. The transversality condition (1.9) is then a direct consequence of Lemma 2.1 and the no-conjugate-points assumption. ∎
Thus, if a thermostat has no conjugate points, then for each $v\in SM$, we can define $G^{*}_{s/u}(v)\subset\Sigma(v)$ by the limiting procedures (1.8). These are the stable and unstable Green bundles. Thanks to the transversality condition (1.9), we see that for each measurable function $p:SM\to\mathbb{R}$ smooth in the direction of the flow, there exist functions $r^{s/u}:SM\rightarrow\mathbb{R}$ so that
$r^{s/u}\beta+\phi_{p}\in G_{s/u}^{*}.$ (2.13)
In general, these functions are measurable. Still, as the next lemma shows, they satisfy a Riccati equation in the flow direction, in which they are always smooth.
Lemma 2.8.
Let $(M,g,\lambda)$ be a thermostat without conjugate points. For each measurable function $p:SM\to\mathbb{R}$ smooth in the direction of the flow, the function $r=r^{s/u}$ characterized by (2.13) satisfies
$r^{2}+(V(\lambda)-2p)r+\kappa_{p}-F(r)=0.$ (2.14)
Proof.
Let us write $G^{*}$ for either $G^{*}_{s}$ or $G^{*}_{u}$. For notational convenience, let us also fix $v\in SM$ and define ${\phi_{p}(t):=\phi_{p}(\varphi_{t}(v))}$, $r(t):=r(\varphi_{t}(v))$, and $\beta(t):=\beta(\varphi_{t}(v))$. If $\eta(t):=d\varphi_{-t}^{-\top}\big{(}r(t)\beta(t)+\phi_{p}(t)\big{)},$ then $\eta(t)\in G^{*}(v)$ for all $t\in\mathbb{R}$. Unraveling the definitions, this means that
$d\varphi_{t}^{-\top}(v)\eta(t)=r(t)\beta(t)+\phi_{p}(t),$
and therefore
$\eta(t)=r(t)\varphi_{t}^{*}\beta(t)+\varphi_{t}^{*}\phi_{p}(t).$
Differentiating with respect to $t$ and setting $t=0$, we can use (2.1) to write
$\dot{\eta}(0)$ $=\dot{r}\beta+r\mathcal{L}_{F}\beta+\mathcal{L}_{F}\phi_{p}$
$=(\dot{r}+rp-\kappa_{p})\beta+(V(\lambda)-p+r)\phi_{p}.$
Since $\eta(0)=r\beta+\phi_{p}$, we obtain
$\dot{\eta}(0)-(V(\lambda)-p+r)\eta(0)=(\dot{r}-\kappa_{p}-rV(\lambda)+2rp-r^{2})\beta.$
The left-hand side belongs to $G^{*}$, and we know that $G^{*}$ is transversal to $\mathbb{H}^{*}$, so the right-hand side must be zero, and the claim follows. ∎
We can now prove Theorem 1.2.
Proof of Theorem 1.2.
One direction is given by Theorem 1.1. In the other direction, suppose the thermostat has no conjugate points. Then, by picking $p=V(\lambda)$ in (2.14), Lemma 2.8 gives us a measurable function $r$ smooth along the flow such that
$r^{2}-V(\lambda)r+\kappa_{0}+F(V(\lambda))-F(r)=0.$
If we now define $p:=V(\lambda)-r$, we get
$\kappa_{p}=\kappa_{0}+F(p)+p(p-V(\lambda))=0,$
as desired. ∎
We also have the tools to prove Theorem 1.3.
Proof of Theorem 1.3.
Recall that $\mathrm{div}_{\mu}F=V(\lambda)$. By Stokes’ theorem, we have
$\int_{SM}F(r)\mu=-\int_{SM}rV(\lambda)\mu.$
Integrating the Riccati equation (2.14), we hence get
$\int_{SM}(r^{2}+2(V(\lambda)-p)r)\mu=-\int_{SM}\kappa_{p}\mu.$
It follows that
$\int_{SM}(\kappa_{p}-(p-V(\lambda))^{2})\mu$ $=-\int_{SM}(r^{2}+2(V(\lambda)-p)r+(p-V(\lambda))^{2})\mu$ (2.15)
$=-\int_{SM}(r+V(\lambda)-p)^{2}\mu\leq 0.$
If the left-hand side is zero, then $r=p-V(\lambda)$ $\mu$-almost everywhere. Now consider
$B_{t}:=\{v\in SM\ |\ r(\varphi_{t}(v))=p(\varphi_{t}(v))-V(\lambda)(\varphi_{t}(v))\},\qquad t\in\mathbb{R}.$
Since $\varphi_{t}$ is smooth, the Radon-Nikodym theorem implies that $B_{t}$ has full measure for every $t\in\mathbb{R}$, hence $B:=B_{0}\cap\big{(}\bigcap_{n\geq 1}B_{1/n}\big{)}$ also has full measure. Thus, we see that for almost every $v\in SM$, we have
$F(r)(v)$ $=\lim_{n\rightarrow\infty}n(r(\varphi_{1/n}(v))-r(v))$
$=\lim_{n\rightarrow\infty}n(p(\varphi_{1/n}(v))-V(\lambda)(\varphi_{1/n}(v))-p(v)+V(\lambda)(v))$
$=F(p-V(\lambda))(v).$
Substituting this into (2.14) yields $\mathbb{K}=0$ $\mu$-almost everywhere. Since $\mathbb{K}$ is a smooth function, we get $\mathbb{K}=0$ everywhere. ∎
In fact, a slight modification of this argument yields the following, which will be useful in the proof of Theorem 1.9.
Lemma 2.9.
Let $(M,g,\lambda)$ be a thermostat without conjugate points. For any finite flow-invariant Borel measure $\nu$ on $SM$, we have
[…]
with equality if and only if $\tilde{\kappa}=0$ $\nu$-almost everywhere.
Proof.
Taking $p=V(\lambda)/2$ and integrating (2.14) with respect to $\nu$, we get
[…]
The left-hand side is zero if and only if $r$ is zero $\nu$-almost everywhere. With the same argument as in the proof of Theorem 1.3, we notice that having $r=0$ $\nu$-almost everywhere implies that $F(r)=0$ $\nu$-almost everywhere. It then follows that $\tilde{\kappa}=0$ $\nu$-almost everywhere. ∎
2.5. Lyapunov exponents
Recall that the Lyapunov exponent at $(v,\xi)\in\Sigma$ is
$\chi(v,\xi):=\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\|d\varphi_{t}^{-\top}(v)\xi\|,$ (2.16)
[…]
$\dot{w}+w^{2}+\tilde{\kappa}=0.$ (2.17)
Furthermore, using (2.7), we get the relationship
$u(t)=w(t)-\frac{1}{2}V(\lambda)(\varphi_{t}(v)).$ (2.18)
We want to relate the exponential growth rate of $\|d\varphi_{t}^{-\top}(v)\xi\|$ to the exponential growth rate of $|y(t)|$ in the case where $\xi\in G_{s/u}^{*}(v)$. For completeness, we recall the following standard Riccati comparison result (see, for example, [Gre54, Lemma 2.1] and the discussion after).
Lemma 2.10.
Fix $\kappa\in\mathcal{C}^{\infty}(\mathbb{R},\mathbb{R})$, and let $w\in\mathcal{C}^{1}(\mathbb{R},\mathbb{R})$ solve the Riccati equation
$\dot{w}+w^{2}+\kappa=0.$
If there is a constant $K>0$ so that $\kappa\geq-K^{2}$, then $|w(t)|\leq K$ for all $t\in\mathbb{R}$.
In particular, compactness and the above lemma imply that globally defined solutions to (2.17) are bounded. This gives us the ingredients to prove the following.
Lemma 2.11.
Let $(M,g,\lambda)$ be a thermostat without conjugate points. If $\xi\in G_{s/u}^{*}(v)$, then there is a constant $C>0$ such that $|x(t)|\leq C|y(t)|$ for all $t\in\mathbb{R}$.
Proof.
Since $\xi\in G_{s/u}^{*}(v)$, we know thanks to the transversality condition in Theorem 1.5 that $y$ never vanishes. Equations (2.3) then imply that
$\frac{x(t)}{y(t)}=(p-V(\lambda))(\varphi_{t}(v))-u(t).$
Since $w$ is bounded, (2.18) along with the fact that $V(\lambda)$ is bounded implies that $u$ is bounded. The claim follows since $p-V(\lambda)$ is also bounded along any orbit. ∎
The following dynamical criterion for the Green bundles will also be useful.
Lemma 2.12.
Let $(M,g,\lambda)$ be a thermostat without conjugate points, and $\xi\in\Sigma(v)$. If $z$ is bounded for all $t\geq 0$ (resp. $t\leq 0$), then $\xi\in G^{*}_{s}(v)$ (resp. $\xi\in G^{*}_{u}(v)$).
Proof.
Let $\xi\in\Sigma(v)$, and suppose $z$ is bounded for all $t\geq 0$. By normalizing, we may assume that $z(0)=1$. Let $z_{t_{0}}$ and $w$ be the solutions to (2.6) satisfying $z_{t_{0}}(0)=1$, $z_{t_{0}}(t_{0})=0$, $w(0)=0$, and $\dot{w}(0)=1$. There must be a family of constants $c_{t_{0}}\in\mathbb{R}$ such that $z=z_{t_{0}}+c_{t_{0}}w$, so it suffices to show that $c_{t_{0}}\rightarrow 0$ as $t_{0}\rightarrow\infty$. To that end, observe that (2.9) implies that, for any $t>0$, we have
$\lim_{t_{0}\rightarrow\infty}c_{t_{0}}=\lim_{t_{0}\rightarrow\infty}\frac{z(t)-z_{t_{0}}(t)}{w(t)}=\frac{z(t)}{w(t)}-\int_{t}^{\infty}\frac{d\tau}{(w(\tau))^{2}}.$
Since $w$ is unbounded by Lemma 2.4 and $z$ is bounded for all $t\geq 0$ by assumption, we get the desired conclusion by taking $t\to\infty$. The same argument with $t_{0}\rightarrow-\infty$ gives us the claim if $z$ is bounded for all $t\leq 0$. ∎
For $\xi\in G_{s/u}^{*}(v)$, we can use Lemma 2.11 and (2.18) to rewrite (2.16) as
[…]
Furthermore, for any $\xi\in G_{s/u}^{*}(v)\setminus\{0\}$, it is clear that $\chi(v,\xi)=\chi(v,\pm\xi/\|\xi\|)$. We write $\chi^{s/u}(v):=\chi(v,\xi)$ for $\xi\in G_{s/u}^{*}(v)\setminus\{0\}$.
Let $\nu$ be a Borel measure on $SM$ which is ergodic for the flow. Oseledets’ theorem [Ose68] says that the limit in (2.16) exists $\nu$-almost everywhere. Furthermore, we have a splitting $\Sigma(v)=E^{0}(v)\oplus E^{-}(v)\oplus E^{+}(v)$ for $\nu$-almost every $v\in SM$, where
$E^{0}(v):=\{\xi\in\Sigma(v)\ |\ \chi(v,\xi)=0\}\qquad\text{ and }\qquad E^{\pm}(v):=\{\xi\in\Sigma(v)\ |\ \mp\chi(v,\xi)<0\}.$
By Lemma 2.12, we have the inclusions
$E^{+}\subseteq G_{u}^{*}\subseteq E^{0}\oplus E^{+}\qquad\text{ and }\qquad E^{-}\subseteq G_{s}^{*}\subseteq E^{0}\oplus E^{-}.$ (2.20)
Note that since we are on a surface and $G_{s/u}^{*}(v)$ are $1$-dimensional subspaces, we either have $E^{+}(v)=G_{u}^{*}(v)$ or $\dim(E^{+}(v))=0$. This leads us to the following.
Lemma 2.13.
Let $(M,g,\lambda)$ be a thermostat without conjugate points, and let $\nu$ be a Borel ergodic measure on $SM$. We have $G_{s}^{*}(v)=G_{u}^{*}(v)$ $\nu$-almost everywhere if and only if $\chi^{u}(v)=\chi^{s}(v)=0$ $\nu$-almost everywhere. Furthermore, if $G_{s}^{*}(v)=G_{u}^{*}(v)$ for all $v\in SM$, then the topological entropy of the thermostat flow is zero.
Proof.
The first statement follows immediately from (2.20). Assume now that we have $G_{s}^{*}(v)=G_{u}^{*}(v)$ for every $v\in SM$. Using Ruelle’s inequality [Rue78], the metric entropy of the flow is zero with respect to any ergodic Borel measure, and hence the metric entropy of any Borel invariant measure is zero using an ergodic decomposition argument (see, for example, [FH19, Theorem 3.3.37]). The variational principle [FH19, Theorem 4.3.7] implies that the topological entropy is zero. ∎
We now show that the Green bundles collapse to a line $\mu$-almost everywhere when the thermostat curvature vanishes everywhere. In the magnetic case, this collapsing happens everywhere, and the converse also holds provided the curvature is non-positive.
Proof of Theorem 1.9.
Suppose $\mathbb{K}=0$. By Theorem 1.1, we know that the thermostat has no conjugate points. Using (2.15) with $p=V(\lambda)$, we get $r^{s/u}=0$ $\mu$-almost everywhere. It follows that $G^{*}_{s}=G^{*}_{u}=\mathbb{R}(\psi_{\lambda}-V(\lambda)\beta)$ $\mu$-almost everywhere.
If we pick $p=V(\lambda)$, then $\kappa_{p}=\mathbb{K}$, and equations (2.3) become
$\begin{cases}\dot{x}+V(\lambda)x=0,\\ \dot{y}+x=0.\end{cases}$
We get the explicit solutions
[…]
In particular, when $V(\lambda)=0$, the solutions are $x(t)=x(0)$ and $y(t)=y(0)-tx(0)$. Therefore, if we start with $x(0)\neq 0$ and $y(0)=0$, which corresponds to a covector in $\mathbb{H}^{*}$, we get
$\lim_{t\to\pm\infty}\dfrac{x(t)}{y(t)}=\lim_{t\pm\infty}-\dfrac{1}{t}=0.$
This tells us that the bundle $d\varphi_{t}^{-\top}(\mathbb{H}^{*}(\varphi_{-t}(v)))$ converges to $\mathbb{R}\psi_{\lambda}$ as $t\to\pm\infty$. This holds everywhere, so we obtain $G^{*}_{s}=G^{*}_{u}=\mathbb{R}\psi_{\lambda}$, proving (a).
Next, observe that it suffices to prove (b) in the setting where $\nu$ is ergodic by taking an ergodic decomposition of the measure. Let $\nu$ be a Borel ergodic measure, and let $\Delta:=\{v\in SM\ |\ \chi^{u}(v)>0\}.$ Whenever the limit exists, let
[…]
By the Birkhoff ergodic theorem, this function is measurable, constant $\nu$-almost everywhere, and
[…]
Thus, if we set $P^{-}:=\left\{v\in SM\ |\ \rho(v)<0\right\}$, then ergodicity implies $\nu(P^{-})=1$ or $\nu(P^{-})=0$. Without loss of generality, we may assume that the limit in (2.19) exists for every $v\in P^{-}$. Once we show that $P^{-}=\Delta$ $\nu$-almost everywhere, (b) will follow from Lemma 2.13. Indeed, if $\tilde{\kappa}(v)<0$ for some $v\in SM$, then Lemma 2.9 along with ergodicity implies that $\nu(P^{-})=\nu(\Delta)=1$, and hence $G_{s}^{*}\neq G_{u}^{*}$ $\nu$-almost everywhere. On the other hand, if $\tilde{\kappa}=0$ $\nu$-almost everywhere, then $\nu(P^{-})=\nu(\Delta)=0$, and hence $G_{s}^{*}=G_{u}^{*}$ $\nu$-almost everywhere.
Integrating (2.17) from 00 to $t$, we get
[…]
In particular, if $v\in P^{-}$ and $w$ corresponds to a covector $\xi\in G_{u}^{*}(v)$, then normalizing the above by $t$, taking the limit, and noting that $w$ is bounded for $t\geq 0$, we have
[…]
Since $\tilde{\kappa}\leq 0$, a comparison argument implies that $w\geq 0$, and a standard analysis lemma implies that $\chi^{u}(v)>0$ (cf. [BP02a, Lemma 3.4.4]). On the other hand, if $v\in\Delta$, then the Cauchy-Schwarz inequality yields
[…]
Thus, $\rho(v)>0$ where the limit exists, and hence $P^{-}=\Delta$ $\nu$-almost everywhere. ∎
3. Transversal Green bundles
The goal of this section is to prove Theorems 1.6 and 1.8. The key property we need to show is that the Green bundles are continuous whenever they are transversal everywhere. For this, we adapt some of the arguments in [Ebe73] to this more general setting, giving a dynamical characterization of the stable and unstable Green bundles. We start with the following corollary of Lemma 2.12.
Corollary 3.1.
Let $(M,g,\lambda)$ be a thermostat with no conjugate points. If $G_{s}^{*}(v)\cap G_{u}^{*}(v)=\{0\}$ for all ${v\in SM}$, then there are no non-trivial bounded solutions to (2.6) for any $(v,\xi)\in\Sigma$.
The next step is to analyze the growth of solutions to (2.6) when they vanish, i.e., when they correspond to vectors in the cohorizontal bundle $\mathbb{H}^{*}$.
Lemma 3.2.
Let $(M,g,\lambda)$ be a thermostat without conjugate points. If $G_{s}^{*}(v)\cap G_{u}^{*}(v)=\{0\}$ for all $v\in SM$, then there exists a constant $A>0$ such that if $z$ is a solution to (2.6) with $z(0)=0$, then $|z(t)|\geq A|z(\tau)|$ for all $t\geq\tau\geq 1$.
Proof.
We argue by contradiction. For each integer $n\geq 1$, pick a non-trivial solution $z_{n}$ to (2.6) and $t_{n}\geq\tau_{n}\geq 1$ such that $z_{n}(0)=0$ and $|z_{n}(t_{n})|\leq(1/n)|z_{n}(\tau_{n})|$. Multiplying by a constant if necessary, we can assume that $\dot{z}_{n}(0)=1$.
For each $n\geq 1$, pick $u_{n}\in[0,t_{n}]$ such that $z_{n}(t)\leq z_{n}(u_{n})$ for all $t\in[0,t_{n}]$. Let $\delta:=\inf_{n\geq 1}u_{n}$. We must have $\delta>0$. If not, then $u_{n}\to 0$ up to picking a subsequence. But then $\lim_{n\to\infty}z_{n}(u_{n})=0$ by compactness of $SM$ and continuity. However, since $t_{n}\geq 1$ for all $n$, we have $z_{n}(u_{n})\geq z_{n}(1)$. Given that $\inf_{n\geq 1}z_{n}(1)>0$ by compactness, this is a contradiction.
Define $w_{n}(t):=z_{n}(t+u_{n})/z_{n}(u_{n})$. Note that each $w_{n}$ satisfies equation (2.6) and
$\begin{cases}w_{n}(-u_{n})=0,\\ w_{n}(0)=1,\\ w_{n}(t_{n}-u_{n})\leq 1/n,\\ w_{n}(t)\leq 1\text{ for }-u_{n}\leq t\leq t_{n}-u_{n}.\end{cases}$
Using Lemma 2.10, we get that there is a uniform constant $K>0$ so that $|\dot{w}_{n}(0)|\leq K$. Moreover, using compactness to work on a convergent subsequence, we have that $\lim_{n\rightarrow\infty}w_{n}=w$. This function $w$ satisfies (2.6) and $w(0)=1$ by continuity, so it is a non-trivial solution. We now argue that we have a contradiction in every possible case:
  1. (1)
    $t_{n}-u_{n}$ and $u_{n}$ both contain bounded subsequences. Up to picking subsequences, we have $t_{n}-u_{n}\to t$ and $u_{n}\to u$. By continuity, we get $w(-u)=0=w(t)=0$, but this is impossible since $-u\leq-\delta<0\leq t$.
  2. (2)
    $u_{n}\to\infty$ and $t_{n}-u_{n}$ contains a bounded subsequence. Up to taking a subsequence, we have $t_{n}-u_{n}\to t$. By continuity, $w(t)=0$ and $w(\tau)\leq 1$ for all $\tau\leq t$. This contradicts Lemma 2.4.
  3. (3)
    $t_{n}-u_{n}\to\infty$ and $u_{n}$ contains a bounded subsequence. Same argument as (2).
  4. (4)
    $t_{n}-u_{n}\to\infty$ and $u_{n}\to\infty$. This implies that $|w(t)|\leq 1$ for all $t\in\mathbb{R}$ by continuity. This contradicts Corollary 3.1. ∎
As mentioned above, thermostats are not reversible. However, reversing the orbit of a thermostat does give rise to the orbit of a thermostat where the intensity is “flipped.” More precisely, let $\mathcal{F}:TM\rightarrow TM$ be the flip map given by $(x,v)\mapsto(x,-v)$, and for $\lambda\in\mathcal{C}^{\infty}(SM)$ define $\lambda^{\mathcal{F}}:=-\lambda\circ\mathcal{F}$. Unraveling the definition, it is clear that if $\gamma$ is a thermostat geodesic for $(M,g,\lambda)$, then the curve $t\mapsto\gamma(-t)$ is a thermostat geodesic for the system $(M,g,\lambda^{\mathcal{F}})$ with initial conditions $(\gamma(0),-\dot{\gamma}(0))$. Furthermore, reversing time and the velocity gives us a correspondence of solutions to (2.6) between the two thermostats. We summarize this observation with the following.
Lemma 3.3.
Let $(M,g,\lambda)$ be a thermostat, and let $\gamma$ be an orbit. The curve $t\mapsto\gamma(-t)$ is a thermostat geodesic for $(M,g,\lambda^{\mathcal{F}})$. Moreover, $z(t)$ is a solution to (2.6) along $t\mapsto\gamma(t)$ for $(M,g,\lambda)$ if and only if $z(-t)$ is a solution to (2.6) along $t\mapsto\gamma(-t)$ for $(M,g,\lambda^{\mathcal{F}})$.
With this, we now have the ingredients to give a dynamical characterization of the Green bundles.
Lemma 3.4.
Let $(M,g,\lambda)$ be a thermostat without conjugate points, and suppose $G_{s}^{*}(v)\cap G_{u}^{*}(v)=\{0\}$ for all $v\in SM$. We have $\xi\in G_{s}^{*}(v)$ (respectively $\xi\in G_{u}^{*}(v)$) if and only if the corresponding solution $z$ to the Jacobi equation (2.6) is bounded for all $t\geq 0$ (respectively $t\leq 0$).
Proof.
Let $\xi\in G_{s}^{*}(v)$ correspond to the solution $z$ to (2.6) with $z(0)=1$, and define $z_{t_{0}}$ as solution to (2.6) with $z_{t_{0}}(0)=1$ and $z_{t_{0}}(t_{0})=0$. If we set $u_{t_{0}}(t):=z_{t_{0}}(-t)$, then $u_{t_{0}}$ is a solution to (2.6) along the orbit $t\mapsto\gamma(-t)$ for the thermostat $(M,g,\lambda^{\mathcal{F}})$. In particular, fixing $t\geq 0$ and considering $t_{0}\geq 1+t$, Lemma 3.2 and Lemma 3.3 imply
$|z_{t_{0}}(t)|=|u_{t_{0}}(-t)|=|u_{t_{0}}(-t_{0}+(t_{0}-t))|\leq\frac{1}{A}.$
By taking $t_{0}\rightarrow\infty$, we get $|z(t)|\leq 1/A$ for all $t\geq 0$. The converse is Lemma 2.12. ∎
Remark 3.5.
Notice that the argument above shows that if $\xi\in G_{s}^{*}(v)$, then the corresponding solution $z$ to (2.6) satisfies $|z(t)|\leq z(0)/A$ for all $t\geq 0$, where $A$ is the uniform constant from Lemma 3.2.
In general, one knows very little about the regularity of the Green bundles. We can now prove that if they are transversal everywhere, they must at least be continuous.
Proposition 3.6.
Let $(M,g,\lambda)$ be a thermostat without conjugate points. If one has $G_{s}^{*}(v)\cap G_{u}^{*}(v)=\{0\}$ for all $v\in SM$, then the Green bundles are continuous.
Proof.
We introduce the notation $z(t;v)$ to emphasize that we are dealing with a solution to (2.6) corresponding to some $\xi\in\Sigma(v)$, and similarly for $m(t;v)$ and $y(t;v)$.
Let $(v_{n})\subseteq SM$ be a sequence such that $v_{n}\rightarrow v$. It suffices to show that $G_{s}^{*}(v_{n})\rightarrow G_{s}^{*}(v)$. To that end, let $z_{n}(t;v_{n})$ be a sequence of solutions to (2.6) such that the corresponding covectors $\xi_{n}$ lie in $G_{s}^{*}(v_{n})$. By normalizing, we may use compactness to assume that $\xi_{n}\rightarrow\xi\in\Sigma(v)$. Further normalize so that $z_{n}(0;v_{n})=1$ for each $n$. Notice that, in coordinates, this implies that $y_{n}(t;v_{n})\rightarrow y(t;v)$ and $\dot{y}_{n}(t;v_{n})\rightarrow\dot{y}(t;v)$ for every $t\in\mathbb{R}$. Let $z(t;v)$ be a solution to (2.6) corresponding to $\xi$. Suppose for contradiction that $z(t;v)$ is not bounded above for all $t\geq 0$. Observe that
$\begin{split}|z(t;v)-z_{n}(t;v_{n})|&\leq\frac{|m(t;v_{n})y(t;v)-m(t;v)y_{n}(t;v_{n})|}{m(t;v)m(t;v_{n})}\\ &\leq\frac{|y(t;v)-y_{n}(t;v_{n})|}{m(t;v_{n})}+\frac{|m(t;v_{n})-m(t;v)||y(t;v)|}{m(t;v)m(t;v_{n})},\end{split}$
and for fixed $t$, the right-hand side tends to zero as $n\to\infty$. By choosing sufficiently large $t$, we have $z_{n}(t;v_{n})\geq 1/A$ for sufficiently large $n$, contradicting Remark 3.5. ∎
Finally, we prove Theorem 1.8.
Proof of Theorem 1.8.
Let $p=0$, and define $w:=-(r^{s}+r^{u})/2$, where $r^{s/u}$ are the functions defined by (2.13). Since $r^{s/u}$ both satisfy the Riccati equation (2.14), we get
$\kappa_{w}$ $=w^{2}-V(\lambda)w+\kappa_{0}+F(w)$
$=\dfrac{\left(r^{s}\right)^{2}}{4}+\dfrac{\left(r^{u}\right)^{2}}{4}+\dfrac{r^{s}r^{u}}{2}-V(\lambda)w+\kappa_{0}+F(w)$
$=-\dfrac{(r^{s})^{2}}{4}-\dfrac{(r^{u})^{2}}{4}+\dfrac{r^{s}r^{u}}{2}$
$=-\dfrac{1}{4}(r^{s}-r^{u})^{2}.$
Since $r^{s}(v)\neq r^{u}(v)$ for all $v\in SM$ by the transversality condition, it follows that $\kappa_{w}<0$. In terms of regularity, the functions $r^{s/u}$ are always smooth along the flow. However, thanks to Proposition 3.6, we also know that they are continuous since the Green bundles are transversal everywhere by assumption. ∎
This allows us to then prove our main result.
Proof of Theorem 1.6.
Given the definitions of $(E^{s})^{*}$ and $(E^{u})^{*}$ in (1.12), it follows that $\Sigma=(E^{s})^{*}\oplus(E^{u})^{*}$. The estimates (1.11) then tell us that $G^{*}_{s}=(E^{s})^{*}$ and $G_{u}^{*}=(E^{u})^{*}$. By Theorem 1.8, we know that there exists a continuous function $p:SM\to\mathbb{R}$, smooth along the flow, such that $\kappa_{p}<0$. We may then apply [MP19, Theorem 3.7]. ∎
4. Examples and counterexamples on the $2$-torus
The following system shows that, when $V(\lambda)\neq 0$, it is possible to have $\mathbb{K}=0$ yet $G^{*}_{s}(v)\cap G^{*}_{u}(v)=\{0\}$ for some $v\in SM$. In particular, we see that Theorem 1.9 (a) is optimal.
Proposition 4.1.
Let $(\mathbb{T}^{2},g)$ be the $2$-torus endowed with a flat Riemannian metric. Define $\lambda\in\mathcal{C}^{\infty}(S\mathbb{T}^{2},\mathbb{R})$ by $\lambda(x,\theta):=\cos(\theta)$. Then $\mathbb{K}=0$ and
$G^{*}_{s}(x,\pi/2)=\mathbb{R}(\psi_{\lambda}-\beta),\qquad$ $G^{*}_{u}(x,\pi/2)=\mathbb{R}\psi_{\lambda},$
$G^{*}_{s}(x,-\pi/2)=\mathbb{R}\psi_{\lambda},\qquad$ $G^{*}_{u}(x,-\pi/2)=\mathbb{R}(\psi_{\lambda}-\beta).$
Proof.
Note that $V(\lambda)(x,\theta)=-\sin(\theta)$. Unraveling the definitions, we get
$\mathbb{K}=\pi^{*}K_{g}-H(\lambda)+\lambda^{2}+F(V(\lambda))=\cos^{2}(\theta)-\cos^{2}(\theta)=0.$
Observe that the sets $\{(x,\theta)\in SM\mid\theta=\pm\pi/2\}$ are invariant by the flow $\varphi_{t}$. When $\theta=\pi/2$, the solutions to (2.21) become
$x(t)=e^{t}x(0),\qquad y(t)=y(0)+x(0)(1-e^{t}).$
Recall that we had set $p=V(\lambda)$, so $p(x,\pi/2)=-1$. Further note that $m(t)=e^{t/2}$. If we start with $x(0)=0$ and $y(0)\neq 0$, which corresponds to a covector in $\mathbb{R}(\psi_{\lambda}-\beta)$, we get that $z(t)=y(0)e^{-t/2}$ is bounded for all $t\geq 0$. By Lemma 2.12, it follows that $G^{*}_{s}(x,\pi/2)=\mathbb{R}(\psi_{\lambda}-\beta)$. On the other hand, if we start with $x(0)\neq 0$ and $y(0)=-x(0)$, which corresponds to a covector in $\mathbb{R}\psi_{\lambda}$, we get that $z(t)=-x(0)e^{t/2}$ is bounded for all $t\leq 0$. By Lemma 2.12, it follows that $G^{*}_{u}(x,\pi/2)=\mathbb{R}\psi_{\lambda}$.
When $\theta=-\pi/2$, we instead have
$x(t)=e^{-t}x(0),\qquad y(t)=y(0)+x(0)(e^{-t}-1),\qquad m(t)=e^{-t/2},$
so we can repeat the same arguments. ∎
To prove Theorem 1.4, we provide the following family of examples.
(a) Dynamics restricted to the fibers of $S\mathbb{T}^{2}$.
(b) Evolution of a single trajectory on $\mathbb{T}^{2}$.
Figure 4. Illustration when $m=2$.
Lemma 4.2.
Let $g$ be a Riemannian metric on $\mathbb{T}^{2}$, and define $\lambda\in\mathcal{C}^{\infty}(S\mathbb{T}^{2},\mathbb{R})$ by
$\lambda(x,\theta):=h(x)+\cos(m\theta),$
where $h\in\mathcal{C}^{\infty}(\mathbb{T}^{2},\mathbb{R})$ and $m$ is an integer $\geq 2$. If
$\|\partial h\|_{\infty}+h^{2}+(m-2)|h|<m-1-K_{g},$
then the thermostat $(\mathbb{T}^{2},g,\lambda)$ has no conjugate points and admits a dominated splitting.
Proof.
Let $p:=V(\lambda)/m$, i.e., $p(x,\theta)=-\sin(m\theta)$, and let us write $\lambda_{0}:=\pi^{*}h$. Then
$\kappa_{p}$ $=\pi^{*}K_{g}-H(\lambda)+\lambda^{2}+F(p)+p(p-V(\lambda))$
$=\pi^{*}K_{g}-H(\lambda_{0})+\lambda_{0}^{2}+(2-m)\lambda_{0}\cos(m\theta)+1-m.$
Therefore, since $m\geq 2$, we obtain
$\kappa_{p}$ $\leq\pi^{*}K_{g}+\|H(\lambda_{0})\|_{\infty}+\lambda_{0}^{2}+(m-2)|\lambda_{0}|+1-m<0.$
The result follows from Theorem 1.1 and [MP19, Theorem 3.7]. ∎
By [PT72], we know that $S\mathbb{T}^{2}$ cannot admit an Anosov flow, and hence this is an example of a thermostat with a dominated splitting which is not Anosov. It is also easy to see that the non-wandering set is $\{(x,\theta)\in S\mathbb{T}^{2}\mid\cos(m\theta)=0,\,\sin(m\theta)=1\}$.